The vapour pressure of a solution of two liquids, A (po = 80 mm, x = 0.4) and B (po = 120 mm, x = 0.6) is found to be 100 mm. It shows that the solution exhibits 


        (A)    Negative deviations from ideal behavior


        (B)    Positive deviations from ideal behavior


        (C)    Ideal behavior


(D)    Positive deviations for lower concentration and negative deviations for higher concentration.


Grade:11

1 Answers

Amit Saxena
10 years ago

It is because ideal PTotal if ideal solution is formed (= 80 × 0.4 + 120 × 0.6 = 32 + 72 = 104 mm)

 

        Since new PTotal < 104 mm \ it shows negative deviation